# LLM-Consensus: Multi-Agent Debate for Visual Misinformation Detection

### v2

@@ Abstract
One of the most challenging forms of misinformation involves the out-of-context (OOC) use of images paired with misleading text, creating false narratives. Existing AI-driven detection systems lack explainability and require expensive finetuning. We address these issues with LLM-Consensus, a multi-agent debate system for OOC misinformation detection. LLM-Consensus introduces a novel multi-agent debate framework where multimodal agents collaborate to assess contextual consistency and request external information to enhance cross-context reasoning and decision-making. Our framework enables explainable detection with state-of-the-art accuracy even without domain-specific fine-tuning. Extensive ablation studies confirm that external retrieval significantly improves detection accuracy, and user studies demonstrate that LLM-Consensus boosts performance for both experts and non-experts. These results position LLM-Consensus as a powerful tool for autonomous and citizen intelligence applications.
misinformation detection,out-of-context image use,LLMs,multimodal models,multi-agent debate,safety
@@ 1 Introduction
Our growing dependence on online channels for news and social networking has been complemented by a surge in exploits of digital misinformation (Aslett et al., 2024; Hasher et al., 1977; Brashier & Marsh, 2020). While many manipulation techniques pose serious threats, one of the most prevalent methods for creating fake online content is the out-of-context (OOC) use of images (pbs, ). This involves using unaltered images in a misleading, false context to convey deceptive information, a strategy that requires minimal technical expertise.
OOC misinformation detection requires a nuanced understanding of the relationship between text and images and the ability to identify when they are misaligned. Detecting these subtle inconsistencies is time-consuming for humans, and (Sultan et al., 2022) shows that time pressure further reduces detection accuracy, exacerbating scalability issues.
As a result, AI-driven tools have gained attention for scaling detection efforts. However, conventional deep learning forensic techniques (Castillo Camacho & Wang, 2021; Heidari et al., 2024; Zhu et al., 2018; Amerini et al., 2021; Hina et al., 2021), designed to detect manipulations like PhotoShop editing (Tolosana et al., 2020; Masood et al., 2023; Farid, 2016; Wang et al., 2019) and AI-generated Deepfakes (mit, ), focus on spotting artifacts from tampering. In contrast, OOC detection requires cross-contextual reasoning, as the deception stems from the misalignment of legitimate images with misleading text.
@@
Figure 1: Overview of LLM-Consensus: Two or more independent agents see the same image-text input and are tasked with detecting whether the input is misinformation or not. After the agents form their independent opinions, they participate in a debate until they converge on the same response or when $n$ debate rounds are completed (whichever is earlier).
Pretrained Large Multimodal Models  (Liu et al., 2024b; OpenAI & et al., 2024; Li et al., 2019; Radford et al., 2021, LMMs) provide a promising direction for detecting OOC use of images for their ability to process both text and image content in tandem. However, using LMMs directly for OOC detection presents several challenges, particularly in the news domain. For instance, news articles often include images that are not directly related to the article’s content. An article about the 2024 U.S. presidential candidates, for example, might feature a close-up of Donald Trump from an unrelated online database. Although the image was taken outside the election period, it is not considered OOC since it doesn’t misrepresent the article’s context. Such cases complicate LMMs’ ability to accurately identify OOC usage based solely on their pre-trained knowledge, as this knowledge may be outdated or insufficiently detailed.
Moreover, even with recent advancements, LMMs are capable of hallucinating and, hence, generating false information (Bai et al., 2024; Liu et al., 2024a). While rapidly improving, they sometimes fail to understand user instructions and intent correctly. We show that off-the-shelf LMMs indeed suffer from these issues, reducing their ability to detect OOC misinformation in practice. While prior work (Qi et al., 2024) has shown that off-the-shelf models can be improved using task-specific fine-tuning, this approach is resource-intensive and requires continual updating to keep up with recent events. Moreover, detecting OOC images only solves part of the problem. The real value lies in being able to explain the OOC use of pictures in human-readable form. It can be instrumental for human validators to observe the model’s line of logic and gain better insight into, and trust in, the classification process.
@@
In this work, we propose a novel LMM-based post-training approach for scalable OOC misinformation detection that simultaneously improves contextual reasoning, provides in-built explainability, and achieves state-of-the-art detection accuracy even without task-specific fine-tuning (see Section 3). Specifically, our framework LLM-Consensus frames the detection problem as a dialectic debate between multiple LMM agents, where, in contrast to prior work (Minsky, 1988; Li et al., 2023a; Du et al., 2023a; Khan et al., 2024)), agents have access to external information retrieval.
Compared to single-agent chain-of-thought approaches (Wei et al., 2024), the use of multiple agents allows for a clean separation of agent contexts, decentralisation of action spaces, and opportunities for parallel computation (Schroeder de Witt et al., 2020; Du et al., 2023b). In addition, due to its compositional nature, both additional human and autonomous agents can be dynamically added to the multi-agent reasoning process, allowing the use of LLM-Consensus as an interactive tool for human experts. To the best of our knowledge, no prior work has used debating LMMs for detecting and explaining OOC image use.
We perform a comprehensive empirical evaluation of our method (see Section 4), including the study of multiple debate configurations. To optimise OpenAI API use, we utilize an experimental pipeline where preliminary experiments are performed using the open-source LLaVA model (Liu et al., 2024b), which is only later replaced with GPT-4o (OpenAI, ) to achieve state-of-the-art performance. We find that LLM-Consensus outperforms both prior work and novel baselines that we introduce, is more robust to various failure modes, and produces coherent explanations that help both human experts and non-experts significantly improve their detection accuracy in user studies. We identify both access to external information retrieval and complete freedom of opinion as key ingredients to the performance of LLM-Consensus. Finally, we discuss current limitations of our method and propose future work toward overcoming the scalability challenges in large-scale online OOC misinformation detection.
@@ 2 Related Work
Recent work has focused on using joint image-text representations to classify an instance as OOC. (Aneja et al., 2022) follow a self-supervised approach to assess whether two captions accompanying an image are contextually similar. They enforce image-text matching during training by formulating a scoring function to align objects in the image with the caption. During inference, they use the semantic similarity between the two captions to classify them as OOC or not. The increased reliance on textual content limits the capabilities of this approach. This work also does not provide explanations for model predictions and is, therefore not interpretable. Moreover, this method works for image caption pairs where captions have information about objects in the image. This is not always the case with news articles (our domain of application), where captions can often just be related to the main content of an article rather than precisely describing the objects in the image. Appendix  A.2 shows an example of the same.
(Abdelnabi et al., 2022) introduce the Consistency Checking Network (CCN), which models aspects of human reasoning across modalities for misinformation detection using external evidence aggregated from the Internet. The CCN employs memory networks to evaluate the consistency of image-caption pairs against retrieved evidence and a CLIP component (Radford et al., 2021) to assess the alignment between the image and caption. The use of external evidence enhances CCN’s classification performance compared to other methods. However, it lacks an explainability component, providing only binary decisions without further explanation.
(Zhang et al., 2024) extend the neural symbolic method (Yi et al., 2019; Zhu et al., 2022) to develop an interpretable cross-modal misinformation detection model that provides supporting evidence for its predictions. They use symbolic graphs based on Abstract Meaning Representation (Banarescu et al., 2013) to detect out-of-context image use. Similarly, (Zhou et al., 2020) introduce Similarity Aware Fake news detection (SAFE), where neural networks jointly learn and analyze features and relationships between text and visual news representations to predict fake news. (Wang et al., 2018) introduce EANN: Event Adversarial Neural Networks to derive event invariant features which can be used to detect fake news that has recently been generated. EANN uses adversarial training to learn multi-modal features independent of news events. These methods require pretraining from scratch and, therefore, don’t benefit from the advanced reasoning capabilities and world knowledge of large pretrained models.
Shalabi et al. (2023) use synthetic multi-modal data to establish the authenticity of image-text pairs. They use BLIP-2 (Li et al., 2023b) to generate a caption for the original image and Stable Diffusion (Rombach et al., 2022) to generate an image for the given original caption. This synthetic data is then used to reason that if the original image and caption are OOC, then the original and generated images should also be OOC as well as the original and generated text. This method relies on synthetic multi-modal data generation, which not only adds an additional computational overhead but also increases dependence on often unreliable synthetically generated data. Therefore, this method can suffer from issues related to generation models, including potential biases that these models may possess. This method also lacks interpretability.
Sniffer (Qi et al., 2024) is the closest to our work. It uses the InstructBLIP (Dai et al., 2023) model to detect OOC image use and provide an explanation for its prediction. It makes use of internal and external knowledge using entity extraction APIs and image-based web searches. Information from all the sources is given to an LLM to predict and explain if an image has been used OOC. Sniffer only uses basic textual information, such as news article titles from websites, to form its external knowledge base. It also requires extensive training to adapt the model to the news domain which adds additional computational overhead and also restricts the generalization abilities of the model to other domains.
@@ 3 Methodology
We present LLM-Consensus, an explainable misinformation detection system that jointly predicts and explains instances of misinformation (Figure 1). Unlike prior work, which largely provides predictions without explanations, our approach uses multiple multi-modal models debating to determine whether an image-text pair constitutes misinformation. We address the question:
@@
Our external retrieval module uses reverse image search to provide agents with real-world context, enhancing their predictions. Using the GPT-4o (OpenAI, ) model, we achieve state-of-the-art performance with detailed, coherent explanations, without requiring domain-specific fine-tuning. This ensures faster generalization to new domains with minimal computational overhead.
@@ 3.1 Debate Modelling
Analogous to real-world conversations, communication between two AI agents can also be structured in a myriad of ways. We explore multiple debating strategies to structure the conversation between agents, all of which are tested and evaluated in our experiments. Instead of simple back-and-forth conversations, we opt for a debating set-up in which agents are asked to frame their own opinions and then defend them to other agent(s). We observe this facilitates more involved and detailed discussions among the models.
Figure 2: Debating Strategies: We experiment with multiple debating strategies. The asynchronous debate setup where agents argue one after the other and take turns presenting their arguments is the best configuration.
@@
One of the core setups we test in our experiments is an asynchronous debating strategy, where models wait for the responses of other participants before generating their own. Figure 2 (a) and (b) illustrate the differences between synchronous and asynchronous debate structures. While synchronous debates, where all participants respond simultaneously, can be faster and more computationally efficient, we find the asynchronous setup more effective. This structure allows models to better identify contextual ambiguities in a more organized and systematic manner, which is crucial for misinformation detection. A key aspect of this setup is the design of model prompts: the debating models are not aware that they are interacting with other AI agents but are instead led to believe they are debating with a human.
@@
We also experiment with an asynchronous debate setup with a judge. Figure 2 (c) shows this setup. In this setup, models participate in an asynchronous debate as usual however, the final decision is made by a judge at the end of the debate. Models are incentivised to structure their arguments in a way that makes them most convincing to the judge. We structure this debate configuration similar to (Khan et al., 2024), where the judge does not have access to all the external information and has to rely only on the debate transcript to decide the final answer.
@@
In this setup, only one agent, the actor, is tasked with deciding whether a given image-text pair is misinformation. The agent generates a response which a skeptic then evaluates. The skeptic is tasked with finding logical errors in the actor’s argument and asking follow-up questions to disambiguate the actor’s response. It is important to note that neither the skeptic nor the actor has access to the ground truth. This setup does not benefit from an ensemble since both models assume different roles and only one agent is tasked with generating the final answer.
Debate with Disambiguation:
Building on the actor-skeptic method, we allow all agents to act as both actors and skeptics. Models generate their own responses and disambiguation queries to refine or challenge other agents’ outputs. These queries are used to retrieve additional information from the Internet, further improving model responses.
@@ 3.2 Prompt Engineering
The debate structure is enabled through prompt engineering. As shown in Figure 1, the first stage involves each AI agent independently determining whether the image-text pair is misinformation, using context from the external retrieval module (details in Appendix A.3). The initial prompt summarizes related news articles and instructs the agent to focus on key image details, like watermarks and flags, to identify inconsistencies.
After forming initial opinions, the agents engage in a debate. The first round uses a unique prompt to help agents adapt to the discussion, while subsequent rounds use a standard prompt incorporating suggestions from previous rounds. Agents must agree or disagree with others’ responses while identifying ambiguities and refining their reasoning.
To prevent blind agreement, the prompt explicitly instructs agents to provide valid reasoning before accepting others’ responses. This promotes stronger, independent stances and enhances the discovery of new information.
@@ 3.3 External Information Retrieval
Figure 3: Structure of the external information retrieval module: We use the Bing Visual Search API (vis, ) to obtain web pages related to a given image, which are then summarised using Llama-13B (Touvron et al., 2023). This summary is then passed to the debating agents as a part of the initial prompt.
A model’s world knowledge is limited by its training data and time frame, but external retrieval allows access to up-to-date information beyond these constraints. Previous work relies on external datasets (Abdelnabi et al., 2022), but these often provide only news article titles, which lack sufficient detail. Since a model’s world knowledge is limited to its training data (and hence a particular time frame), incorporating external retrieval allows the model to access information beyond this training data (and time frame). Previous work makes use of pre-existing external retrieval-based datasets (Abdelnabi et al., 2022) to supplement external information related to an image-caption pair. However, we find this information lacking in detail since it is limited to the title of a news article. Full access to the article content can significantly enhance an agent’s ability to assess whether an image-caption pair constitutes misinformation. To address this, we propose a dedicated external information retrieval module, which improves accuracy when integrated into the pipeline. The module operates in two stages:
3.3.1 API-Based Information Retrieval
The Bing Visual Search API is used for the task of obtaining web pages related to a given image (vis, ) . A given image from the dataset is used to obtain a list of web pages completely and partially related to the image. We take the top three matching web pages in which the image appears. We believe these web pages contain sufficient information to allow the agent to develop a general understanding of the context in which the image is originally used. Since the community-accepted dataset for this task; NewsCLIPpings (Luo et al., 2021), contains images from articles published more than ten years ago, some images do not result in any web pages or viable search results. In such a scenario, we simply do not pass any external context to the agent and only rely on the agent’s existing knowledge base. Since this is relevant for only a very small number of examples in the dataset, it does not adversely affect system performance.
3.3.2 Summarization using LLM
Once the top three web pages are identified, we scrape their text to gather contextual information about the image. This textual data is often too lengthy for direct input to the agent, so we use the Llama-13B (Touvron et al., 2023) model to generate concise summaries, focusing on key details to help the agent better understand the external context. For non-English web pages, the LLM struggles with summarization, so we add a filter to exclude them. Although this restricts the system to English, it does not impact performance due to the dataset’s predominantly English-language content. Multi-lingual support could be added by translating text to English before summarization.
3.4 Coherent Reasoning
All the different components of LLM-Consensus are brought together in this stage of the pipeline. Each multi-modal agent is employed to participate in the best-debating set-up with the relevant prompts and is asked to detect a given image-text pair as misinformation and provide an explanation for the same. The agents also have access to external information related to the image through the external retrieval module. The final decision of the system is obtained once the debate terminates, which is after a certain number of debate rounds or after all agents converge to a common response, whichever is earlier.
@@ 4.1 Dataset
We perform a series of experiments and report results on the NewsCLIPpings dataset (Luo et al., 2021). The dataset is built based on the VisualNews (Liu et al., 2020) dataset, which consists of image-caption pairs from four news agencies: BBC, USA Today, The Guardian, and The Washington Post. The NewsCLIPpings dataset is created by generating OOC samples by replacing an image in one image-caption pair with a semantically related image from a different image-caption pair. CLIP (Radford et al., 2021) is used to retrieve semantically similar images for a given caption. We report results on the Merged-Balanced version of the dataset, which has balanced proportions of all the retrieval strategies and positive/negative samples. The training, validation and test sets have 71,072, 7,024 and 7,264 samples, respectively. We use the NewsCLIPpings dataset for evaluating our model’s efficacy in identifying out-of-context images and information due to its well-established use in prior work, its relative scale compared to other available datasets, and the lack of more recent large-scale benchmarks in this domain. Though no single dataset can capture the full diversity of real-world misinformation, NewsCLIPpings provides a standardized evaluation setting that facilitates comparability with existing methods.

### v3

# MAD-Sherlock: Multi-Agent Debate for Visual Misinformation Detection

@@ Abstract
One of the most challenging forms of misinformation involves pairing images with misleading text to create false narratives. Existing AI-driven detection systems often require domain-specific finetuning, limiting generalizability, and offer little insight into their decisions, hindering trust and adoption. We introduce MAD-Sherlock, a multi-agent debate system for out-of-context misinformation detection. MAD-Sherlock frames detection as a multi-agent debate, reflecting the diverse and conflicting discourse found online. Multimodal agents collaborate to assess contextual consistency and retrieve external information to support cross-context reasoning. Our framework is domain- and time-agnostic—requiring no finetuning–yet achieves state-of-the-art accuracy with in-depth explanations. Evaluated on NewsCLIPpings, VERITE, and MMFakeBench, it outperforms prior methods by 2%, 3%, and 5%, respectively. Ablation and user studies show that the debate and resultant explanations significantly improve detection performance and improves trust for both experts and non-experts, positioning MAD-Sherlock as a robust tool for autonomous citizen intelligence.
multi-agent debate,agents,misinformation
@@ 1 Introduction
The rise of online news and social media has been paralleled by a surge in digital misinformation [4, 14, 6]. Among the most widespread tactics is out-of-context (OOC) image use [13], where unaltered images are paired with misleading text to deceive, requiring little technical skill.
OOC detection demands nuanced reasoning to identify misalignment between images and accompanying text. This task is time-consuming for humans, and detection accuracy drops under time pressure [37], limiting scalability.
AI tools offer a path forward, but traditional forensic methods [8, 15, 46, 2, 16] focus on tampering artifacts (e.g., Photoshop edits [12, 40] or Deepfakes [38, 26, 36]) and are ill-suited for OOC detection, which hinges on cross-contextual reasoning rather than visual manipulation.
@@
Figure 1: Overview of MAD-Sherlock: Two or more independent agents see the same image-text input and are tasked with detecting whether the input is misinformation or not. After the agents form their independent opinions, they participate in a debate until they converge on the same response or when $n$ debate rounds are completed (whichever is earlier).
Pretrained large multimodal models (LMMs) offer a promising foundation for detecting OOC image use by jointly interpreting text and images [23, 30, 19, 34]. However, applying them to news content is challenging. News articles often feature loosely related but contextually appropriate images—e.g., a pre-2024 photo of Donald Trump used in election coverage—which makes OOC detection difficult using pretraining alone. LMMs also suffer from hallucinations, misinterpret intent [5, 22], and lack up-to-date knowledge. Off-the-shelf models exhibit these limitations, reducing their reliability. Fine-tuning helps [33], but is costly and requires frequent updates. Crucially, beyond detection, models must provide clear, human-readable explanations to support understanding and trust in their decisions.
@@
In this work, we propose a novel post-training framework for scalable OOC misinformation detection that improves contextual reasoning, offers built-in explainability, and achieves state-of-the-art accuracy without task-specific fine-tuning (see Section 3). Our method frames the detection problem as a dialectic debate between LMM agents, augmented with external information retrieval.
Unlike single-agent chain-of-thought methods [42], our multi-agent setup enables context separation, decentralized reasoning, and parallelization [35, 11]. Prior work [27] shows agents struggle to maintain diverse perspectives within a single context window; our approach addresses this via modular, compositional interaction. It also allows seamless integration of human or autonomous agents, making MAD-Sherlock a flexible tool for expert oversight. To our knowledge, this is the first use of multi-agent LMM debate for both detecting and explaining OOC image use.
MAD-Sherlock is backbone-agnostic, compatible with open- and closed-source models. We prototype and tune with LLaVA [23] to reduce API costs, then deploy GPT-4o [29] for final results. By avoiding task-specific fine-tuning, our method remains broadly applicable across domains and time periods.
We evaluate on three benchmarks—NewsCLIPpings, VERITE, and MMFakeBench—and achieve new state-of-the-art performance across all. MAD-Sherlock outperforms prior methods and baselines, demonstrates robustness, and produces detailed, interpretable explanations. In user studies, these explanations significantly improve detection accuracy for both experts and non-experts.
Key contributors to performance include external retrieval and agent independence. We discuss limitations and outline future directions toward scalable, general-purpose AI for public good.
@@ 2 Related Work
Recent work has explored joint image-text representations for OOC classification. Aneja et al. [3] use a self-supervised approach to enforce image-text alignment via an object-caption matching score, classifying OOC instances by caption similarity. However, the method is heavily text-dependent and lacks interpretability.
Abdelnabi et al. [1] introduce the Consistency Checking Network (CCN), which leverages memory networks and CLIP for image-caption consistency against external evidence, improving accuracy but offering only binary outputs without explanations.
Zhang et al. [43] employ AMR-based symbolic graphs for interpretable OOC detection with evidence. Similarly, Zhou et al. [44] propose SAFE for joint text-visual reasoning in fake news detection. Wang et al. [41] present EANN, which uses adversarial training to extract event-invariant multimodal features. These methods require training from scratch and lack the advanced reasoning and knowledge of large pretrained models.
Sniffer [33] is most similar to our work. It uses InstructBLIP [9] to detect OOC image use and generate explanations by aggregating internal and external knowledge from entity extraction APIs and image-based web searches. However, its external knowledge is limited to basic textual content (e.g., article titles), and adapting it to the news domain requires extensive training, reducing generalizability and increasing computational overhead.
The CRAVE framework [10] clusters retrieved multimodal evidence into coherent narratives, then uses an LLM to produce fact-checking judgments with natural language explanations. While effective, treating all clusters equally may amplify fringe perspectives and distort the evidence base.
Liu et al. [24] introduce MMD-Agent, a single-agent framework that hierarchically decomposes misinformation detection into textual, visual, and cross-modal subtasks before reasoning over their outputs to make a final prediction.
@@ 3 Methodology
We present MAD-Sherlock, an explainable misinformation detection system that jointly predicts and explains instances of misinformation (Figure 1). Unlike prior work, which largely provides predictions without explanations, our approach uses multiple multi-modal models debating to determine whether an image-text pair constitutes misinformation. We address the question:
@@
Our external retrieval module uses reverse image search to provide agents with real-world context, enhancing their predictions. Using the GPT-4o [29] model, we achieve state-of-the-art performance with detailed, coherent explanations, without requiring domain-specific fine-tuning. This ensures faster generalization to new domains with minimal computational overhead.
@@ 3.1 Debate Modelling
Analogous to real-world conversations, communication between two AI agents can also be structured in a myriad of ways. We explore multiple debating strategies to structure the conversation between agents, all of which are tested and evaluated in our experiments and informed by the work of Brown-Cohen et al. [7].
Figure 2: Debating Strategies: We experiment with multiple debating strategies to evaluate which performs best on our task.
@@
We experiment with an asynchronous debating strategy, where models wait for others’ responses before generating their own (Figure 2a vs. b). While synchronous debates are faster and more efficient, the asynchronous setup proves more effective for identifying contextual ambiguities—critical in misinformation detection. Notably, models are prompted to believe they are debating a human rather than another AI.
@@
We also explore an asynchronous debate setup with a judge (Figure 2c), where models debate as usual, but a final decision is made by a judge based solely on the debate transcript. Following Khan et al. [18], the judge lacks access to external information, incentivizing models to structure arguments that are maximally persuasive.
@@
In this setup, a single actor determines whether an image-text pair constitutes misinformation. A skeptic then critiques the actor’s reasoning, probing for logical flaws and ambiguities. Neither agent has access to the ground truth, and since their roles are distinct, this configuration does not benefit from ensembling.
@@ 3.2 Prompt Engineering
The debate is structured via prompt engineering (Figure 1). In the first stage, each agent independently assesses whether the image-text pair is misinformation, using context from an external retrieval module (see Appendix A.3). Prompts summarize relevant articles and emphasize visual cues (e.g., watermarks, flags) to guide initial judgments. Agents then debate: the first round uses a tailored prompt to initiate discussion; later rounds use a standard prompt incorporating prior responses. Agents must agree or disagree with peers while refining reasoning, and prompts explicitly discourage blind agreement by requiring justification for any alignment.
@@ 3.3 External Information Retrieval
Figure 3: Structure of the external information retrieval module: We use the Bing Visual Search API to obtain web pages related to a given image, which are then summarised using Llama-13B [39]. This summary is then passed to the debating agents as a part of the initial prompt.
A model’s world knowledge is bounded by its training data and time frame, but external retrieval enables access to up-to-date context. Prior work leverages retrieval-based datasets [1], though these typically include only news headlines, which are often too sparse for reliable inference. Full article access can significantly improve misinformation detection. To this end, we introduce a two-stage external retrieval module that enhances accuracy when integrated into the pipeline:
API-Based Information Retrieval
We use the Bing Visual Search API to retrieve web pages related to each image. For each image, we collect the top three matching pages in which it appears, assuming these provide sufficient context for understanding the image’s original use. Since NewsCLIPpings [25] includes articles over a decade old, some images yield no search results. In these rare cases, we omit external context and rely solely on the agent’s prior knowledge, which has minimal impact on overall performance.
Summarization using LLM
After retrieving the top three web pages, we scrape their text and use LLaMA-13B [39] to generate concise summaries focused on key contextual details. Since the model struggles with non-English content, we filter out non-English pages–limiting the system to English but without affecting performance, as the dataset is predominantly English. Multilingual support could be added via translation prior to summarization.
Coherent Reasoning
This stage of the pipeline integrates all components of MAD-Sherlock. Each multimodal agent engages in the optimal debate setup using relevant prompts and is tasked with determining and explaining whether a given image-text pair constitutes misinformation. Agents also access contextual information via the external retrieval module. The system’s final decision is produced when the debate concludes, which is either after a fixed number of rounds or once all agents converge on a shared response, whichever occurs first.
@@ 4.1 Dataset
We conduct experiments on three datasets: NewsCLIPpings, VERITE, and MMFakeBench [25, 32, 24]. This selection balances well-established benchmarks like NewsCLIPpings with newer, more diverse datasets such as VERITE and MMFakeBench, providing both continuity with prior work and coverage of recent advances.
The NewsCLIPpings dataset builds on VisualNews [21], which contains image-caption pairs from BBC, USA Today, The Guardian, and The Washington Post. OOC samples are created by replacing an image with a semantically related one from a different pair. We use the Merged-Balanced version, with 71,072 training, 7,024 validation, and 7,264 test samples, offering balanced retrieval strategies and label distributions. Its scale and prior adoption make it well-suited for evaluating our model’s ability to detect out-of-context content.
We also evaluate on the VERITE dataset [32], a recent benchmark for multimodal misinformation detection that controls for unimodal bias. Each image-caption pair is constructed so that neither modality alone determines veracity. VERITE includes three subsets—All, True vs. OOC, and True vs. MC—each testing different aspects of multimodal reasoning. Captions and images appear in both truthful and misleading contexts, promoting balanced, joint reasoning. The test split contains 338 true, 324 OOC, and 338 MC samples. We report results on the combined misinformation set (MC + OOC), with disaggregated results in Appendix A.4.
We also evaluate on MMFakeBench [24], a benchmark for multimodal misinformation detection in mixed-source scenarios. The validation set includes 1,000 image-caption pairs spanning three major categories—textual, visual, and cross-modal distortions—each with 12 subtypes, capturing the complexity of real-world misinformation.